### v1

# Tyche: Stochastic In-Context Learning for Medical Image Segmentation

Abstract
Existing learning-based solutions to medical image segmentation have two important shortcomings. First, for most new segmentation task, a new model has to be trained or fine-tuned. This requires extensive resources and machine-learning expertise, and is therefore often infeasible for medical researchers and clinicians. Second, most existing segmentation methods produce a single deterministic segmentation mask for a given image. In practice however, there is often considerable uncertainty about what constitutes the correct segmentation, and different expert annotators will often segment the same image differently. We tackle both of these problems with Tyche, a model that uses a context set to generate stochastic predictions for previously unseen tasks without the need to retrain. Tyche differs from other in-context segmentation methods in two important ways. (1) We introduce a novel convolution block architecture that enables interactions among predictions. (2) We introduce in-context test-time augmentation, a new mechanism to provide prediction stochasticity. When combined with appropriate model design and loss functions, Tyche can predict a set of plausible diverse segmentation candidates for new or unseen medical images and segmentation tasks without the need to retrain. Code available at: https://github.com/mariannerakic/tyche/.
Figure 1: Tyche: the first in-context stochastic segmentation framework. Human annotators (top) can handle a wide variety of tasks, and different annotators often produce differing segmentations. Existing automated methods (middle) are typically task-specific and provide only one segmentation per image. Tyche (bottom) can capture the disagreement among annotators across many modalities and anatomies without retraining or fine-tuning.
1 Introduction
Segmentation is a core step in medical image analysis, for both research and clinical applications. However, current approaches to medical image segmentation fall short in two key areas. First, segmentation typically involves training a new model for each new modality and biomedical domain, which quickly becomes infeasible given the resources and expertise available in biomedical research and clinical environments. Second, models most often provide a single solution, whereas in many cases, the target image contains ambiguous regions, and there isn’t a single correct segmentation. This ambiguity can arise from noisy or low contrast images, variation in the task definition, or human raters’ interpretations and downstream goals [12, 55]. Failure to take this ambiguity into account can affect downstream analysis, diagnosis, and treatment.
Recent work tackles these issues separately. In-context learning methods generalize to unseen medical image segmentation tasks, employing an input context or prompt to guide inference [19, 129, 128]. These methods are deterministic and predict a single segmentation for a given input image and task.
Separately, stochastic or probabilistic segmentation methods output multiple plausible segmentations at inference, reflecting the task uncertainty [11, 67, 95]. Each such model is trained for a specific task, and can only output multiple plausible segmentations at inference for that task. Training or fine-tuning a model for a new task requires technical expertise and computational resources that are often unavailable in biomedical settings.
We present Tyche, a framework for stochastic in-context medical image segmentation (Figure 1). Tyche includes two variants for different settings. The first, Tyche-TS (Train-time Stochasticity), is a system explicitly designed to produce multiple candidate segmentations. The second, Tyche-IS (Inference-time Stochasticity), is a test time solution that leverages a pretrained deterministic in-context model.
Tyche takes as input the image to be segmented (target), and a context set of image-segmentation pairs that defines the task. This enables the model to perform unseen segmentation tasks upon deployment, omitting the need to train new models. Tyche-TS learns a distribution of possible label maps, and predicts a set of plausible stochastic segmentations. Tyche-TS encourages diverse predictions by enabling the internal representations of the different predictions to interact with each other through a novel convolutional mechanism, a carefully chosen loss function and noise as an additional input. In Tyche-IS, we show that applying test time augmentation to both the target and context set in combination with a trained in-context model leads to competitive segmentation candidates.
We make the following contributions.
  • We present the first solution for probabilistic segmentation for in-context learning. We develop two variants to our framework: Tyche-TS that is trained to maximize the quality of the best prediction, and Tyche-IS, that can be used straightaway with an existing in-context model.
  • For Tyche-TS, we introduce a new mechanism, SetBlock, to encourage diverse segmentation candidates. Simpler than existing stochastic methods, Tyche-TS predicts all the segmentation candidates in a single forward pass.
  • Through rigorous experiments and ablations on a set of twenty unseen medical imaging tasks, we show that both frameworks produce solutions that outperform existing in-context and interactive segmentation benchmarks, and can match the performance of specialized stochastic networks trained on specific datasets.
2 Related Work
Biomedical segmentation is a widely-studied problem, with recent methods dominated by UNet-like architectures [107, 6, 51]. These models tackle a wide variety of tasks, such as different anatomical regions, different structures to segment within a region, different image modalities, and different image settings. With most methods, a new model has to be trained or fine-tuned for each combination of these. Additionally, most models don’t take into account image ambiguity, and provide a single deterministic output.
Multiple Predictions. Uncertainty estimation can help users decide how much faith to put in a segmentation [26] and guide downstream tasks. Uncertainty is often categorized into aleatoric, uncertainty in the data, and epistemic, uncertainty in the model [29, 61]. In this work, we focus on aleatoric uncertainty. Medical images are also heteroscedastic in that the degree of uncertainty varies across the image.
Different strategies exist to capture uncertainty. One can assign a probability to each pixel [62, 45, 52, 76], or use contour strategies and difference loss functions to predict the largest and smallest plausible segmentations [73, 132]. These strategies however do not capture the correlations across pixels. To address this, some methods generate multiple plausible label maps given an image [95, 11, 67, 68, 130]. To achieve this, one can directly model pixel correlations, such as through a multivariate Gaussian distribution (with low rank) covariance [95], or more complex distributions [15]. Alternatively, various frameworks combine potentially hierarchical representations for UNet-like architectures with variational auto-encoders [67, 11, 68]. More recently, diffusion models have been used for ensembling [130] or to produce stochastic segmentations [105, 133]. Some methods explicitly model the different annotators to capture ambiguity [48, 100, 111, 124]. But these methods do not apply to our framework where the number of annotators and their characteristics are unknown.
Most of the models above involve sophisticated modeling or lengthy runtimes, and need to be trained on each segmentation task. In Tyche, we build on intuition across these methods, but combine a more efficient mechanism with an in-context strategy to predict segmentation candidates.
In-context Learning. Few-Shot frameworks use a small set of examples to generalize to new tasks [31, 80, 98, 117, 101, 113, 134], sometimes by fine-tuning an existing pretrained model [32, 99, 125, 120]. In-context learning segmentation methods (ICL) use a small set of examples directly as input to infer label maps for a task [19, 9, 63, 129, 129, 63]. This enables them to generalize to new tasks. For example, UniverSeg uses an enhanced UNet-based architecture to generalize to medical image segmentation tasks unseen during training [19]. We build on these ideas to enable segmentation of new tasks without the need to re-train, but expand this paradigm to model stochastic segmentations.
Test Time Augmentation. The test-time augmentation (TTA) strategy uses perturbations of a test input and ensembles the resulting predictions. Existing TTA frameworks model accuracy [116, 33, 119, 64], robustness [25], and estimates of uncertainty [90, 5]. Test-time augmentation has been applied to diverse anatomies and modalities including brain MRI and retinal fundus [3, 5, 49, 97, 53, 127]. Prior work has formalized the variance of a model’s predictions over a set of input transformations as capturing aleatoric uncertainty [126, 127, 5].
Tyche’s use of TTA is distinct from prior work. Instead of ensembling segmentations over perturbations of a test input or pixel-wise estimates of uncertainty, Tyche extends TTA to the in-context setting and uses the individual TTA predictions to model uncertainty.
Figure 2: Tyche Model Schematic. The target $x^{t}$, context set ${(x^{t}_{j},y^{t}_{j})}_{j=1}^{S}$, and noise images $\{z_{k}\}_{k=1}^{K}$ are inputs to the network. The architecture employs UNet-like levels, but uses SetBlocks that enable interactions between the context set and the target segmentation candidates.
Figure 3: CrossBlock Mechanism The CrossBlock involves interactions between a single feature and a set of features and outputs new feature for the target and new features for each.
Figure 4: SetBlock Mechanism. SetBlock enables interactions between the set of features from the context set and the set of features from the prediction candidates. It outputs two sets of features, one for the context and one for the prediction candidates.
3 Method
For segmentation task $t$, let $\{(x^{t}_{j},y^{t}_{j})\}_{j=1}^{N}$ be a dataset with images $x^{t}$ and label maps $y^{t}$. Typical segmentation models learn a different function $\hat{y}^{t}=g_{\theta^{t}}(x^{t})$ with parameters $\theta^{t}$ for each task $t$, where $\hat{y}^{t}$ is a single segmentation map prediction.
We design Tyche as an in-context learning (ICL) model using a single function for all tasks:
$\hat{y}_{k}^{t}=f_{\theta}(x^{t},z_{k},\mathcal{S}^{t}).$ (1)
This function, with global parameters $\theta$, captures a distribution of label maps $\{\hat{y}^{t}_{k}\}_{k=1}^{K}$, given target $x^{t}$, context set $\mathcal{S}^{t}=\{x^{t}_{j},y^{t}_{j}\}_{j=1}^{S}$ defining task t, and noise $z_{k}\sim\mathcal{N}(\bm{0},\mathbb{I})$. We use this modelling strategy in two ways: we either explicitly train a network to approximate the model $f_{\theta}(\cdot)$ in Tyche-TS, or design a test-time strategy to approximate $f_{\theta}(\cdot)$ using an existing (pretrained) deterministic in-context network in Tyche-IS.
3.1 Tyche-TS
In Tyche-TS, we explicitly train a neural network for $f_{\theta}(\cdot)$ that can make different predictions given the same image input $x^{t}$ but different noise channels $z_{k}$. We model interaction between predictions, and employ a loss that encourages diverse solutions (Figure 2).
3.1.1 Neural Network
We use a convolutional architecture focused on interacting representations of sets of flexible sizes using a modified version of the usual UNet structure [107].
Inputs. Tyche-TS takes as input the target $x^{t}$, a set of K Gaussian noise channels $z_{k}$, and a context set, $\mathcal{S}^{t}$.
Layers. Each level of the UNet takes as input a set of K candidate representations and S context representations. We design each level to encourage communication between the intermediate elements of the sets, and between the two features of the segmentation candidates. The size $K$ is flexible and can vary with iterations.
SetBlock. We introduce a new operation called SetBlock, which interacts the candidate representations $U=\{u_{i}\}_{i=1}^{K}$, with the context representations $V=\{v_{i}\}_{i=1}^{S}$, illustrated in Figure 4. We use the CrossBlock [19] as a building block for this new layer. The $\text{CrossBlock}(u,V)\rightarrow(u^{\prime},V^{\prime})$ compares an image representation $u$ to a context set representation $V$ through convolutional and averaging operations, and outputs a new image representation $u^{\prime}$ and a new set representation $V^{\prime}$ (Figure 3). $\text{SetBlock}(U,V)\rightarrow(U^{\prime},V^{\prime})$ builds on CrossBlock and performs a set to set interaction of the entries of $U$ and $V$:
$\bar{u}$ $=1/m{\textstyle\sum}_{i=1}^{m}u_{i}$ (2)
$\bar{u}^{\prime},V^{\prime}$ $=\text{CrossBlock}(\bar{u},V)$ (3)
$u_{i}^{\prime}$ $=\text{Conv}_{m}\left(u_{i}||\bar{u}^{\prime}\right),\quad i=1,\ldots,K\,$ (4)
$u_{i}^{\prime}$ $=\text{Conv}_{u}\left(u_{i}^{\prime}\right),\quad i=1,\ldots,\,K$ (5)
$v_{i}^{\prime}$ $=\text{Conv}_{v}\left(v_{i}\right),\quad i=1,\ldots,\,S,$ (6)
[…]
3.1.2 Best candidate Loss
Typical loss function compute the loss of a single prediction relative to a single target, but Tyche-TS produces multiple predictions and has one or more corresponding label maps.We optimize
$\small\mathcal{L}(\theta;\mathcal{T})=\mathbb{E}_{t\in\mathcal{T}}\left[\mathbb{E}_{(x^{t},y^{t}_{r}),\mathcal{S}^{t}}\left[\mathcal{L}_{seg}\left(\{\hat{y}_{k}\},y^{t}_{r}\right)\right]\right],$ (7)
with
$\small\mathcal{L}_{seg}(\{\hat{y}_{k}\},y)=\,\min_{k}\;\mathcal{L}_{Dice}\left(y_{k},y\right),$ (8)
where $y_{r}^{t}$ is a segmentation from rater $r$, and $\mathcal{L}_{Dice}$ is a weighted sum of soft Dice loss [94] and categorical crossentropy. By only back-propagating through the best prediction among $K$ candidates, the network is encouraged to produce diverse solutions [22, 41, 66, 81].
3.1.3 Training Data
We employ a large dataset of single- and multi-rater segmentations across diverse biomedical domains. We then use data augmentation [19], as described in B.3.
We add synthetic multi-annotator data by modelling an image as the average of four blobs representing four raters (Figure 10). Each blob is white disk $b_{i}$ deformed by a random smoothed deformation field $\phi_{i}$. The synthetic image is a noisy weighted sum of raters: $\sum_{i=1}^{4}w_{i}(b_{i}\circ\phi_{i})$ where $\circ$ represents the spacial warp operation.
3.1.4 Implementation Details
We use a UNet-like architecture of 4 SetBlock layers for the encoder and decoder, with 64 features each and Leaky ReLU as activation function. We use the Adam optimizer and a learning rate of 0.0001. At training, we have a fixed number of candidates per sample $K_{\text{tr}}=8$. At inference, we consider different numbers of candidates.
3.2 Tyche-IS
In Tyche-IS, we first train (or use an existing trained) deterministic in-context segmentation system:
$\hat{y}^{t}=h_{\theta}(x^{t},\mathcal{S}^{t}).$
We then introduce a test-time in-context augmentation strategy to provide stochastic predictions:
$\hat{y}^{t}_{k}$ $=f_{\theta}(x^{t},z_{k},\mathcal{S}^{t})$ (9)
$=h_{\theta}(aug(x^{t},z_{k},\mathcal{S}^{t})),$ (10)
where $\tilde{x}^{t},\tilde{\mathcal{S}}^{t}=aug(x^{t},z_{k},\mathcal{S}^{t})$ is an augmentation function.
3.2.1 Augmentation Strategy
Test time augmentation for single task networks $y=g_{t}(x)$ applies different transforms to an input image $x$:
$\small\tilde{x}_{k}=a_{\phi}(x,z_{k}),$ (11)
where $\phi$ are augmentation parameters and $z_{k}$ is a random vector. A final prediction is then obtained by combining several predictions of augmented images. Most commonly, the combining function averages the predictions:
$y=\frac{1}{k}\sum_{k}g_{t}(\tilde{x}_{k}),$ (12)
where the sum operates pixel-wise.
We introduce in-context test-time augmentation (ICTTA) prediction as another mechanism to generate diverse stochastic predictions.
We apply augmentation to both the test target $x^{t}$ and the context set $\mathcal{S}^{t}$:
$\small(\tilde{x_{i}}^{t},y^{t}_{i})=(a_{\phi}(x^{t}_{i}),y^{t})$ (13)
$\tilde{\mathcal{S}}^{t}=\{a_{\phi}\left(x_{j}^{t}\right),y_{j}^{t}\}_{j=1}^{S}.$ (14)
We repeat this process $K_{i}$ times to obtain $K_{i}$ stochastic predictions:
$\hat{y}_{k}=f_{\theta}(\tilde{x_{i}}^{t},z_{k},\tilde{\mathcal{S}^{t}})$ (15)
We only apply intensity based transforms, to avoid the need to invert the segmentations back. We apply Gaussian noise, blurring and pixel intensity inversion. We detail the specific augmentations in B.2.
Figure 5: Visualization of predictions for three different samples, 1 per row. Left: LIDC-IDRI. Right: Hippocampus dataset. The leftmost columns are raters’ annotations. The 4 last columns are model predictions. Tyche provides a set of prediction that is diverse and matches the raters, for tasks unseen at training time.
Figure 6: Single annotator visualization for different models. We show three example images that show very different corresponding segmentation. Tyche can output plausible segmentation for single annotator data with varying degrees of variability in the segmentation. Methods with an asterisk are upper baselines.
4 Experimental Setup
4.1 Data
We evaluate our method using a large collection of biomedical and synthetic datasets. Most datasets include a single manual segmentation for each example, while a few have several raters per image.
Data Splits. We partition each dataset into development, validation, and test splits. We assign each dataset to an in-distribution set (I.D.) or an out-of-distribution set (O.D.). We train exclusively on the development splits of the I.D. datasets, and use the validation splits of the I.D. datasets to tune parameters. We use the validation splits of the O.D. datasets for final model selection. We report results on the test splits of the O.D. datasets.
For each use case, we sample the context from each dataset’s corresponding development set. As a result, the network doesn’t see any of the O.D. datasets at training time.
We distinguish between single annotator data and multi-annotator data.
Single-Annotator Data. For single annotator data, we build on MegaMedical used in recent publications [19, 131] and employ a collection of 73 datasets, of different public biomedical domains and different modalities [14, 1, 56, 77, 39, 59, 79, 86, 103, 122, 110, 115, 104, 138, 84, 139, 83, 19, 35, 137, 21, 34, 24, 16, 17, 42, 54, 57, 74, 75, 82, 78, 85, 118, 87, 121, 106, 69, 96, 7, 36, 43, 71, 72, 88, 89, 91, 112, 40, 50, 2, 135]. MegaMedical spans a variety of anatomies and modalities, including brain MRI, cardiac ultrasound, thoracic CT and dental X-ray. We also use synthetic data involving simulated shapes, intensities, and image artifacts [19, 44]. The single-annotator datasets used for out-of-domain (O.D.) testing are: PanDental [1], WBC [139], SCD [104], ACDC [14], and SpineWeb [138].
Multi-Annotator Data. For multi-annotator I.D. data, we use four datasets from Qubiq [92]: Brain Growth, Brain Lesions, Pancreas Lesions, and Kidney. We also simulate a multi-rater dataset consisting of random shapes (blobs). For the O.D. multi-annotator data we use four datasets. One contains hippocampus segmentation maps on brain MRIs from a large hospital. We crop the volumes around the hippocampus [67] to focus on the areas where the raters disagree. The second is a publicly available lung nodule dataset, LIDC-IDRI [4]. This dataset is notable for the substantial inter-rater variability. It contains 1018 thoracic CT scans, each annotated by 4 annotators from a pool of 12 annotators. Finally, we also use retinal fundus images, STARE [47], annotated by 2 raters, and prostate data from the MICCAI 2021 QUBIQ challenge [92], annotated by 6 raters on two tasks. Single and multi-annotator combined, our O.D. group contains 20 tasks unseen at training time (some datasets have several tasks).
4.2 Evaluation
We evaluate our method by analysing individual prediction quality and distribution of predictions, both qualitatively and quantitatively. We also examine model choices through an ablation study.
A main use case of stochastic segmentation is to propose a small set of segmentations to a human rater, who can select the most appropriate one for their purpose. For this scenario, a model can be viewed as good if at least one prediction matches what the rater is looking for. We thus employ the best candidate Dice metric.
In the multi-annotator setting, we evaluate using both best candidate Dice score, also called maximum Dice score, as well as Generalized Energy Distance (GED)[13, 109, 123]. GED is commonly used in the stochastic segmentation literature to asses the difference between the distribution of predictions and the distribution of annotations [67, 11, 95, 105, 133]. GED has limitations, such as rewarding excessive prediction diversity [105]. Let $\mathcal{Y}$ and $\hat{\mathcal{Y}}$ be the set of annotations, GED is defined as:
$\small D^{2}_{GED}(\mathcal{Y},\hat{\mathcal{Y}})=2\mathbb{E}\left[d(p,\hat{p})\right]-\mathbb{E}\left[d(p,p^{\prime})\right]-\mathbb{E}\left[d(\hat{p},\hat{p}^{\prime})\right],$ (16)
where $p,p^{\prime}\sim\mathcal{Y}$, $\hat{p},\hat{p}^{\prime}\sim\hat{\mathcal{Y}}$ and $d(\cdot,\cdot)$ is a distance metric. We use the Dice score [30].
4.3 Benchmarks
Tyche is the first method to produce stochastic segmentation predictions in-context. Consequently, we compare Tyche to existing benchmarks, each of which achieves only a subset of our goals.
In-Context Stochastic Automatic
SENet
UniverSeg
SegGPT
Prob. UNet
PhiSeg
CIMD
SAM-based
Tyche
Table 1: Summary of evaluated methods used and their properties. Only Tyche is both stochastic and in-context, and does not require user interaction.
In-Context Methods. We compare to deterministic frameworks that can leverage a context set: a few-shot method, SENet [108], and two in-context learning (ICL) methods, UniverSeg [19] and SegGPT [129]. We train UniverSeg and SENet with the same data split strategies and the same sets of augmentation transforms as for Tyche. For SegGPT, we use the public model, trained on a mix of natural and medical images. Figure 13 in the Supplemental Material shows that UniverSeg trained with additional data outperforms its public version.
Stochastic Upper Bounds. We compare to task-specialized probabilistic segmentation methods that are trained-on and perform well on specific datasets. We independently train Probabilistic UNet [67], PhiSeg[11] and CIDM, a recent diffusion network [105], on each of the 20 held-out tasks. For each task, we train three model variants: no augmentation, weak augmentation, and as much augmentation as for the Tyche targets. For each benchmark variant, we train on a O.D. development split and select the model that performs best on the corresponding O.D. validation split. We then compare these benchmarks to Tyche on the held-out O.D. test splits.
These models are explicitly optimized for the datasets on which they are evaluated, unlike Tyche, which does not use those datasets for training. Since these are trained, tuned and evaluated on the O.D. datasets splits, something we explicitly aim to avoid in the problem set up as it is not easily done in many medical settings, they serve as upper bounds on performance.
Interactive Segmentation Methods. We compare to two interactive methods: SAM [66] and SAM-Med2D [23]. These methods can provide multiple segmentations, but, unlike Tyche, require human interaction, which is outside our scope. SAM also has a functionality to segment all elements in an image, however less optimized for medical imaging. We assume that the SAM-based models have access to the same information as the ICL methods: several image-segmentation pairs as context to guide the segmentation task. We fine-tune SAM using our I.D. development datasets. To replace the human interaction, we provide a bounding box, the average context label map, and 10 clicks, 5 positive and 5 negative as input. With SAM-Med2D, we use a bounding box, and several positive and negative clicks as input. For both SAM and SAM-Med2D, we generate clicks and bounding box from the average context label map.
We use one iteration of interaction, and sample different plausible segmentation candidates by sampling different sets of clicks and different averaged context sets.
Table 1 summarizes the features of all the methods. Additional information on the benchmarks is provided in the Supplemental Material.
$GED^{2}$ ($\downarrow$) Hippocampus LIDC-IDRI Prostate Task 1 Prostate Task 2 STARE
SAM $0.57\pm 0.02$ $0.90\pm 0.01$ $0.20\pm 0.03$ $0.31\pm 0.06$ $0.89\pm 0.06$
Interactive SAM-Med2d $0.93\pm 0.02$ $1.01\pm 0.01$ $0.80\pm 0.09$ $0.78\pm 0.11$ $1.52\pm 0.05$
Tyche-IS $\bm{0.21\pm 0.01}$ $0.41\pm 0.01$ $0.12\pm 0.02$ $0.20\pm 0.05$ $0.73\pm 0.03$
I-C & Stochastic (Ours) Tyche-TS $0.22\pm 0.01$ $\bm{0.40\pm 0.01}$ $\bm{0.09\pm 0.02}$ $\bm{0.15\pm 0.03}$ $\bm{0.62\pm 0.03}$
PhiSeg $0.14\pm 0.01$ $0.33\pm 0.01$ $0.12\pm 0.01$ $0.17\pm 0.05$ $1.22\pm 0.02$
ProbaUNet $0.13\pm 0.01$ $0.51\pm 0.01$ $0.08\pm 0.01$ $0.18\pm 0.05$ $0.76\pm 0.06$
Stochastic Upper Bound CIDM $0.17\pm 0.01$ $0.42\pm 0.01$ $0.14\pm 0.02$ $0.26\pm 0.04$ $0.87\pm 0.05$
Table 2: Generalized Energy Distance for different models with a context size of 16 for in-context methods and a number of predictions set to 8. Lower is better. Tyche outperforms interactive and in-context baselines, and matches stochastic upper bounds.
Max Dice ($\uparrow$) Hippocampus LIDC-IDRI Prostate Task 1 Prostate Task 2 STARE
UniverSeg $0.84\pm 0.01$ $0.67\pm 0.01$ $0.91\pm 0.01$ $0.88\pm 0.03$ $0.51\pm 0.02$
SegGPT $0.10\pm 0.01$ $0.68\pm 0.01$ $0.94\pm 0.01$ $0.89\pm 0.03$ $0.02\pm 0.01$
In-Context SENet $0.68\pm 0.01$ $00.00\pm 0$ 0.00 $0.83\pm 0.02$ $0.83\pm 0.02$ $0.30\pm 0.03$
SAM $0.71\pm 0.01$ $0.55\pm 0.01$ $0.90\pm 0.01$ $0.85\pm 0.03$ $0.50\pm 0.03$
Interactive SAM-Med2d $0.52\pm 0.01$ $0.42\pm 0.01$ $0.62\pm 0.04$ $0.64\pm 0.06$ $0.21\pm 0.03$
Tyche-IS $0.87\pm 0.01$ $0.90\pm 0.00$ $0.94\pm 0.01$ $0.91\pm 0.01$ $0.52\pm 0.03$
I-C & Stochastic (Ours) Tyche-TS $\bm{0.88\pm 0.01}$ $\bm{0.91\pm 0.00}$ $\bm{0.95\pm 0.01}$ $\bm{0.93\pm 0.01}$ $\bm{0.60\pm 0.02}$
PhiSeg $0.88\pm 0.00$ $0.91\pm 0.00$ $0.93\pm 0.01$ $0.91\pm 0.02$ $0.15\pm 0.01$
ProbaUNet $0.91\pm 0.00$ $0.86\pm 0.01$ $0.95\pm 0.00$ $0.91\pm 0.03$ $0.59\pm 0.02$
Stochastic Upper Bound CIMD $0.84\pm 0.01$ $0.92\pm 0.00$ $0.93\pm 0.01$ $0.87\pm 0.02$ $0.41\pm 0.04$
Table 3: Best candidate Dice score for different models with a context size of 16 for in-context methods and a number of predictions set to 8. Higher is better. Tyche outperforms interactive and in-context baselines, and matches stochastic upper bounds.
4.4 Experiments
We evaluate all models on the multi-annotator and single-annotator O.D. data. We then analyze the Tyche variants individually and perform an ablation study on each to validate parameter choices. Finally, we compare the GPU inference runtimes and model parameters.
In the Supplemental Material, we analyze further the noise given as input, the context set, the number of predictions, the SetBlock and the candidate loss. We also provide additional performance metrics and per dataset results. We also compare the performance of Tyche and PhiSeg in a few-shot setting. Finally, we provide additional visualizations.
Inference Setting. We use a fixed context of 16 image-segmentation pairs, because existing in-context learning systems show minimal improvements beyond this size [19]. Because there is variability in performance depending on the context sampled, we sample 5 different context sets for each datapoint and average performance. Similarly, for the stochastic upper bounds and interactive methods, we do 5 rounds of sampling $K_{i}$ samples.
5 Results
5.1 Comparison to Benchmarks
Multi-Annotator O.D. Data. We evaluate on the datasets where multiple annotations exist for each sample. Figure 5 shows that, for both the lung nodules and the hippocampus datasets, Tyche predictions are diverse and capture rater diversity, even though these datasets are out-of-domain. Tables 2 and 3 show that both versions of Tyche outperform the interactive and deterministic benchmarks on all datasets except for Prostate Task 1, on which SegGPT has similar performance. Using a paired Student t-test, we find that Tyche-TS outperforms Tyche-IS in terms of maximum Dice score, with $p<10^{-10}$. We find no statistical difference between the two methods in terms of GED.
Figure 7: Best candidate Dice Score for single annotator data aggregated per task. Tyche outperforms the in-context and interactive segmentation benchmarks, and approaches the stochastic upper bounds. Error bars represent the 95% confidence interval.
Single-Annotator Data. Figure 6 shows examples of predictions for Tyche and the corresponding benchmarks for the single-annotator datasets. Tyche produces a more diverse set of candidates than its competitors. Figure 7 compares all plausible models in terms of aggregate best candidate Dice score, except for CIDM, which underperformed for single-annotator data with a mean best candidate Dice score of: $0.673\pm 0.032$. For clarity, we only present the full Figure in the Supplemental Material. Tyche performs better than the deterministic and interactive frameworks, and similarly to Probabilistic UNet, one of the upper bound benchmarks that is trained on the O.D. data. A paired Student t-test shows that Tyche-IS produces statistically higher GED ($p=0.044$) than Tyche-TS, but we find no statistical difference in terms of best candidate Dice. We hypothesize that Tyche-IS is competitive because of the implicit annotator characterization provided by the context.
5.2 Tyche Analysis
We analyze Tyche variants and study the influence of two important parameters: the number of inference-time candidate predictions $K_{i}$ and the size of the context set $\|\mathcal{S}\|$.
Influence of the number of prediction $K_{i}$. We study how the number of predictions impacts the best candidate Dice score, keeping the context size constant. Figure 8 shows that for Tyche-TS, the best candidate Dice score rises with the number of predictions, but with diminishing returns.
Figure 8: Best candidate Dice Score as the number of candidate prediction increases. The largest improvements are usually obtained for a small number of predictions. The error bars represent the 95% confidence interval.
Influence of context size $\|\mathcal{S}\|$. Figure 9 shows that Tyche-TS is capable of leveraging the increased context size to improve its best candidate Dice score, and that a context size of 16 is sufficient to achieve most of the gain.
Figure 9: Best candidate Dice Score per dataset as context size increases. A context size of 16 is already large enough to obtain a reasonable best candidate Dice. The error bars represent the 95% confidence interval.
Blob SetBlock Std Max. DSC($\uparrow$) $GED^{2}$($\downarrow$)
$0.810\pm 0.01$ $0.349\pm 0.04$
$0.771\pm 0.02$ $0.425\pm 0.05$
$0.802\pm 0.01$ $0.425\pm 0.05$
$\bm{0.811\pm 0.01}$ $\bm{0.298\pm 0.03}$
Target CS CS+ Max. DSC($\uparrow$) $GED^{2}$ ($\downarrow$)
$0.776\pm 0.02$ $0.477\pm 0.04$
$0.700\pm 0.02$ $0.410\pm 0.05$
$0.561\pm 0.02$ $0.867\pm 0.05$
$\bm{0.813\pm 0.01}$ $\bm{0.333\pm 0.04}$
$0.808\pm 0.01$ $0.358\pm 0.04$
Table 4: Ablation Study for Tyche variants. Top: Tyche-TS, without simulated multi-annotator data, with SetBlock, with Standard Deviation in SetBlock. Bottom: Tyche TeS, with Target, Context and Large Context augmentations.
Ablation. Table 4 illustrates several ablations on Tyche design choices. For Tyche-TS, we evaluate the following variants: no simulated multi-annotator images, no SetBlock and finally, using the standard deviation of candidate feature representations in addition to the mean in the SetBlock (“Std”). We compare the models using best candidate Dice averaged across tasks.
Table 4 shows that the simulated multi-annotator data provides negligible improvement, as does adding the standard deviation. However, SetBlock is a crucial part to improve the best candidate Dice score.
We study performances for three types of TTA in Tyche-IS: on the target, on the context (CS), and on the context while also including the non-augmented context (CS+): $(S,\mathcal{G}(S))$. Table 4 shows that adding noise to only one of the target and context yields sub-optimal performance, while augmenting both the target and context improves performances.
5.3 Inference Runtime
We compare the inference runtime by predicting 8 segmentation candidates with each method, and repeat the process 300 times. We use an NVIDIA V100 GPU. Table 5 shows that Tyche is significantly faster and smaller than SegGPT and CIDM, yet, not as fast as some task-specific stochastic models. Tyche-IS has fewer parameters than Tyche-TS, but needs additional inference time.
Inference Time (ms) Parameters
UniverSeg $96.62\pm 0.61$ $1.2$M
SegGPT $2857.19\pm 4.38$ $370$M
SENet $14.91\pm 0.21$ $0.89$M
FT-SAM $1036.75\pm 4.61$ $94$M
SAM-Med2D $188.8\pm 7.58$ $91$M
PhiSeg $11.35\pm 0.672$ $21.1$M
ProbaUNet $8.44\pm 0.46$ $5$M
CIDM $1.7\times 10^{5}\pm 2748$ $85.6$M
Tyche-IS $128.57\pm 2.626$ $1.2$M
Tyche-TS $18.09\pm 0.61$ $1.7$M
Table 5: Inference Runtime and Model Parameters for 8 predictions and a context size of 16.
6 Conclusion
We introduced Tyche, the first framework for stochastic in-context segmentation. For any (new) segmentation task, Tyche can directly produce diverse segmentation candidates, from which practitioners can select the most suitable one, and draw a better understanding of the underlying uncertainty. Tyche can generalize to images from data unseen at training and outperforms in-context and interactive benchmarks. In addition, Tyche often matches stochastic models on tasks for which those models have been specifically trained. Tyche has two variants, one designed to optimize the best segmentation candidate, with fast inference time, and a test-time augmentation variant that can be used in combination with existing in-context learning methods. We are excited to further study the different types of uncertainty captured by Tyche-TS and Tyche-IS. We will also extend the capabilities of Tyche with more complex support sets, including variable annotators and multiple image modalities.
7 Acknowledgement
Research reported in this paper was supported by the National Institute of Biomedical Imaging and Bioengineering of the National Institutes of Health under award number R01EB033773. This work was also supported in part by funding from the Eric and Wendy Schmidt Center at the Broad Institute of MIT and Harvard as well as Quanta Computer Inc. Finally, some of the computation resources required for this research was performed on computational hardware generously provided by the Massachusetts Life Sciences Center.
Supplementary Material
Appendix A Overview
We first present a brief overview of the analysis and information in this Supplemental Material.
Tyche Model Choices.
We present additional data details, including for MegaMedical, multi-annotator data, and simulated data. We give additional details on the Tyche-TS training strategy, as well as the augmentations used at inference for Tyche-IS. We also detail how we trained each benchmark, and their respective limitations.
We provide an intuition behind the best candidate Dice loss, and show that with Tyche, it is possible to optimize for objectives that apply to the candidate predictions as a group, and show results when optimizing GED.
Tyche Analysis.
For Tyche-TS, we investigate four aspects: noise, context set, number of predictions, and SetBlock. We show how different noise levels impact the predictions. Moreover, we give examples of how prediction changes when the context or the noise changes. We also show that the number of predictions at inference impacts the diversity of the segmentation candidates predicted.
For Tyche-IS, we investigate how different augmentations affect performance.
We also analyze a scenario where only few of annotated examples are available, using the LIDC-IDRI dataset. We compare Tyche with these samples in the context to PhiSeg, trained on these few annotated samples.
Further Evaluation.
Given the nature of stochastic segmentation tasks, no single metric fits all purposes, and the optimal metric depends on the downstream goals. We provide sample diversity and Hungarian Matching, showing that Tyche performs well on these as well.
We also present performance per dataset and show that the relative performance of each model stays generally unchanged, even though the difficulty of each dataset is widely different.
We provide additional visualizations on how Tyche and the baselines perform for each datasets, both on single and multi-annotator data.
Data split. All analysis results in this supplemental material use the validation set of the out-of-distribution datasets, to avoid making modeling decisions on the test sets.
Appendix B Tyche Model Choices
We present the implementation details for the data, our Tyche models, and the benchmarks.
B.1 Medical Image Data
Megamedical. We build on the dataset collection proposed in [19, 131], using the similar preprocessing methods. The complete list of datasets is presented Table 10.
Processing. Each image is normalized between 0 and 1 and is resized to $128\times 128$. When the full 3D volume is available, we take two slices for each task: the slice in the middle of the volume and the slice with the largest count of pixels labelled for that task.
Task Definition. We consider a task as labelling a certain structure from a certain modality for a certain dataset. If a given medical image has different structures labeled, we consider each structure a different task. For 3D volumes, we consider each axis a different task.
Synthetic Data. We use a set of synthetic tasks to enhance the performance of our networks, similar to [19]. Some examples are shown in Figure 11.
Synthetic Multi-Annotator Data. We use synthetic data to encourage diversity in our predictions. Figure 10 shows examples of blob targets with the corresponding simulated annotations.
Figure 10: Synthetic Multi-Annotated Examples. Each blob is generated by deforming a white disk using a random smooth deformation field, each representing a different rater. The blobs are then averaged to form the target image to segment.
Figure 11: Synthetic Data Examples. Example of training images on the single annotator synthetic data. Each row is an example of a target-context pair corresponding to a different task.
B.2 Training of Tyche-TS
We train Tyche-TS with a context size of 16 and a batch size of 4. We use a learning rate of 0.0001, Adam as an optimizer, and a kernel size of 3. We apply the augmentations shown Table 6. To obtain $K$ segmentation candidates, we generate $K$ noise samples $z_{k}\sim\mathcal{N}(0,\mathbb{I})$. We duplicate the input $K$ times and concatenate each noise sample with a duplicated target to form $K$ stochastic inputs to the network.
Augmentations $p$ Parameters
degrees $\in[-25,25]$
translate$\in[0,0.1]$
Random Affine 0.25 scale$\in[0.9,1.1]$
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast$\in[0.5,1.5]$
$\alpha\in[1,2.5]$
Elastic Transform 0.8 $\sigma\in[7,9]$
Sharpness 0.25 sharpness$=5$
Flip Intensities 0.5 None
$\sigma\in[0.1,1.0]$
Gaussian Blur 0.25 k=5
$\mu\in[0,0.05]$
Gaussian Noise 0.25 $\sigma\in[0,0.05]$
(a) In-Task Augmentation
Augmentations $p$ Parameters
degrees $\in[0,360]$
translate$\in[0,0.2]$
Random Affine 0.5 scale$\in[0.8,1.1]$
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast$\in[0.8,1.2]$
$\sigma\in[0.1,1.1]$
Gaussian Blur 0.5 $k=5$
$\mu\in[0,0.05]$
Gaussian Noise 0.5 $\sigma\in[0,0.05]$
$\alpha\in[1,2]$
Elastic Transform 0.5 $\sigma\in[6,8]$
Sharpness 0.5 sharpness$=5$
Horizontal Flip 0.5 None
Vertical Flip 0.5 None
Sobel Edges Label 0.5 None
(b) Task Augmentation
Table 6: Set of augmentations used to train Tyche-TS. We distinguish between augmentations aimed at increasing the diversity inside a task (Top) and the augmentations aimed at increasing the diversity of tasks (Bottom). An augmentation is applied with probability $p$ (second column).
Figure 12: Example Augmentations for Tyche-IS. We show for samples from different datasets how three new segmentation candidates are generated. Starting from the top row, augmentations are applied to both the target and context set to obtain new prediction. Each row represents a new target-context pair. We do not show the corresponding labels as for Tyche-IS, we only augment with intensity-based transforms.
B.3 Network for Tyche-IS
For the Tyche-IS network, we use the baseline UniverSeg [19], trained with the same data as Tyche-TS, same batch size, and same context size. At test time, we use the augmentations in Table 7. Figure 12 shows example augmentation on different samples.
Tyche-IS Augment. $p$ Parameters
$\sigma\in[0.1,1.0]$
Gaussian Blur 0.25 $k=5$
$\mu\in[0,0.05]$
Gaussian Noise 0.25 $\sigma\in[0,0.05]$
Flip Intensities 0.5 None
Sharpness 0.25 sharpness=5
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.25 contrast$\in[0.5,1.5]$
Table 7: Augmentations used for Tyche-IS. We focus on intensity transforms, to avoid inverting the prediction. For each image, an augmentation is sampled with probability $p$.
Figure 13: Improvement of UniverSeg by training with more data. By training UniverSeg on more data, we obtain a better average Dice score than the official UniverSeg release.
B.4 Benchmarks
We distinguish three types of benchmarks: in-context methods, that can take as input a context set, the interactive frameworks, and the task-trained specialized upper bounds.
In-context baselines. Because they were trained on medical data, we use SegGPT and SAM-Med2D as provided in their official release. We train from scratch UniverSeg and SENet. We use a batch size of 4 and the same set of data augmentation transforms as the ones used for Tyche-TS to encourage within task and across task diversity. Training UniverSeg on the same datasets as Tyche improves performances compared to the official UniverSeg release, as shown in Figure 13.
Specialized benchmarks. For the specialized models (CIDM, PhiSeg, and Probabilistic UNet), we train a model for each task, for a total of 20 tasks. For each task, we train three model variants, one for each different data augmentation scheme, shown Table 8. We select the model with the data augmentation strategy that does best on the out-of-distribution validation set. We use a batch size of 4. For CIDM and Probabilistic UNet, we use the official PyTorch release. We found Probabilistic UNet particularly unstable to train on some datasets when applying augmentations. To avoid diverging loss values, we had to initialize the Probabilistic UNet models from versions trained on datasets without augmentation. For PhiSeg, the official TensorFLow release was unstable, and we worked with the authors to run a more recent PyTorch version of their code. Per the authors’ request, we smooth the predicted segmentations of CIDM to remove irregularities in the final segmentations.
Augmentation Set 1 $p$ Parameters
None 1 None
(a) No Augmentation. The images are given to the specialized model as is.
Light Augmentation $p$ Parameters
$\sigma\in[0.1,1.5]$
Gaussian Blur 0.5 $k=7$
$\mu\in[0,0.1]$
Gaussian Noise 0.5 $\sigma\in[0,0.1]$
$\alpha\in[1,2]$
Variable Elastic Transform 0.25 $\sigma\in[6,8]$
(b) Light Augmentation. Gaussian bur, Gaussian noise, elastic transforms are each applied to the target with probability $p$.
Heavy Augmentation $p$ Parameters
$\sigma\in[0.1,1.5]$
Gaussian Blur 0.5 $k=7$
$\mu\in[0,0.1]$
Gaussian Noise 0.25 $\sigma\in[0,0.1]$
$\alpha\in[1,2]$
Elastic Transform 0.25 $\sigma\in[6,8]$
degrees $\in[0,360]$
translate$\in[0,0.2]$
Random Affine 0.5 scale$\in[0.8,1.1]$
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast$\in[0.5,1.5]$
Horizontal Flip 0.5 None
Vertical Flip 0.5 None
Sharpness 0.5 sharpness$=5$
(c) Full Set of Augmentations. To form the last augmentation set, we add the following transforms: Affine transform, Sharpness, Brightness and Flips.
Table 8: Set of Augmentations used to train the specialized benchmarks. For evaluation, we select the model with the training augmentations that does best on the out-of-distribution validation set.
Interactive Methods. We use the SAM-Med2D and SAM as interactive segmentation baselines. In our setting, user interaction is not available. We simulate it by averaging the ground truth labels from the context set and sample clicks and bounding boxes from averages.
We use SAM-Med2D’s official release as it is an adaptation of SAM exclusively meant for 2D medical image segmentation tasks. We found empirically that SAM-Med2D performs best when 3 positive and 2 negative clicks are sampled. For the negative clicks, we sample areas inside the bounding box that are not covered by the context set.
We also fine-tune SAM on our data. We sample 5 positive clicks and 5 negative clicks uniformly. We also provide a bounding box and the average of the context label maps as input.
Appendix C Advantages of candidates loss
C.1 Intuition behind best candidate loss
The best candidate loss only evaluates the candidate that yields the lowest loss. We provide intuition for this loss function (Figure 14).
Given as input a target $x^{t}$ and a context $\mathcal{S}^{t}$, the model outputs $K$ segmentation candidates $\hat{y}_{j}$. Despite potentially high target ambiguity, for each training iteration we use only one annotation $y^{t}_{r}$, .
The loss element $l(\hat{y}_{k},y^{t}_{r})$ captures volume overlap between a candidate prediction $\hat{y}_{k}$ and the rater annotation $y^{t}_{r}$. If a regular loss is used across all predictions, then each prediction tends towards the lowest expect cost over the space of possible segmentations for that image, leading to a mean segmentation among raters. This is particularly harmful when target ambiguity is high (high rater disagreement). Then, the mean segmentation may be very different than any one rater and may not be representative of the ambiguity. For the best candidate loss, the model is encouraged to make very different guesses for different segmentation candidates. This increases the chance that one prediction matches the rater being used as ground truth at that iteration.
Figure 14: Best candidate Loss Explanation. Because of ambiguity in the target, there are multiple plausible ground truths. With a standard Dice loss, all the candidates segmentation converge to the mean segmentation. The best candidate loss enables the model to explore different plausible segmentation, since even one proposal matching the rater segmentation used at that iteration leads to a good loss value.
Figure 15: Tyche-TS with GED loss (yellow) improved test-time GED performance (lower is better).
C.2 Distribution loss function
Since Tyche-TS enables multiple candidates that can coordinate with one another, we can employ loss functions that apply to a collection of raters and predictions. We demonstrate this concept by fine-tuning Tyche-TS, using a $GED^{2}$ loss function.
Figure 15 shows that the corresponding model produces samples with lower $GED^{2}$, as expected.
Figure 16: Prediction variability as a function of injected noise. Examples of predictions for three subjects with three different types of input noise. Top to Bottom: zero noise, Gaussian noise constant across segmentation candidates, and randomly sampled Gaussian noise. The context set is fixed. With random noise as an input Tyche can output diverse segmentation candidates.
Figure 17: Prediction variability as a function of the context set. The context set significantly contributes to the variability of the output. For three different subjects, we show how the corresponding segmentation candidates are affected by different context sets. The random Gaussian noise given as input is fixed for each set of candidates.
Figure 18: Tyche-TS with no noise gives similar results to UniverSeg. Setting the noise to 0 for Tyche-TS gives similar best candidate Dice score and similar $GED^{2}$ to UniverSeg. Top: Generalized Energy Distance. Bottom: Best candidate Dice Score.
Figure 19: Setting a number of predictions larger than 1 allows for more diversity. The outputs of Tyche-TS are a lot more uniform when the number of prediction is set to 1 than when it is set to 5.
Figure 20: Removing the SetBlock leads to less diversity in the output. The segmentation candidates output by Tyche-TS are more diverse than the candidates of a Tyche-TS model trained without the SetBlock interaction.
Appendix D Tyche Analysis
We first analyse the different components that lead to variability in Tyche-TS: the noise, the context set, the number of prediction in on forward pass, and the SetBlock mechanism. We then investigate how the size of the context set and the number of predictions impact performances. For Tyche-IS, we investigate different plausible sets of augmentations to apply. Finally, we show that Tyche can produce good results even in data scarcity settings when a stochastic methods trained solely on a few samples would have limited performance.
D.1 Diversity from noise
Figure 16 shows how three predictions vary depending on three noise variants: zero-noise, constant noise across candidates, and variable noise. Both zero noise and constant noise lead systematically to identical segmentation candidates. For each example of segmentation candidates, the context set is fixed.
Figure 18 shows that setting the noise to 0 for Tyche-TS produces similar performances to UniverSeg, both for GED and best candidate Dice score.
D.2 Diversity from context
Figure 17 shows example predictions for three subjects and three context sets. Different contexts lead to different predictions, even if the noise is the same, sometimes drastically. For instance, the predictions for subject 1 and context 3 contain a candidate with no annotations.
D.3 Diversity in the number of predictions
While Tyche-IS predicts each segmentation sample sequentially, Tyche-TS has a one-shot mechanism that predicts all the candidates at once. One may wonder how the number of predictions requested at inference impacts the output of Tyche-TS. Figure 19 shows that with the number of prediction set to 1, Tyche-TS loses a lot of its diversity compared to 5 predictions.
This is also shown quantitatively in Figure 21, where Generalized Energy Distance decreases as the number of predictions increases.
D.4 Diversity from the SetBlock
Tyche-TS has an intrinsic mechanism to encourage the segmentation candidates to be diverse: SetBlock. Figure 20 visually compares the segmentations predicted with the mechanisms and the segmentations predicted by a Tyche model trained without the SetBlock. Without the SetBlock, the predictions output by Tyche are a lot less diverse.
D.5 Size of the Context Set
Generally, a larger context set improves performances. Figure 22 shows that Generalized Energy Distance improves as the size of the context set increases. However, the improvement decreases beyond 16 samples.
Figure 21: $GED^{2}$ for Tyche as a function of the number of predictions, $K_{i}$. Performances improve with the number of prediction but with diminishing returns.
Figure 22: $GED^{2}$ for Tyche as a function of the inference context size. Performances improve as the context size increases but with diminishing returns.
D.6 Augmentations
We study different augmentation for Tyche-IS and how the influence the quality of predictions. We consider four settings: no augmentation, only light augmentation such as Gaussian Noise and Gaussian Blur, described Table 7, the Tyche-IS augmentations shown Table 7 and the Tyche-IS with slightly stronger parameters. shown Table 9. The results are shown in Figure 23 both aggregated across datasets and for each dataset individually. Overall, adding augmentations improves the best candidate Dice score. However, it can degrades the quality of the predictions, for instance for STARE. The augmentation we selected for Tyche-IS is the most promising so far, without requiring inversion of the transform applied.
Figure 23: Performance of different augmentation schemes for Tyche-IS. We consider four augmentations: none, light, the one used in Tyche-IS and a stronger version of the later. Top: Per Dataset. Bottom: Overall. The one that we selected for Tyche-IS is the most promising.
Tyche-IS High Aug. $p$ Parameters
$\sigma\in[0.5,1.0]$
Gaussian Blur 0.25 $k=5$
$\mu\in[0.4,0.5]$
Gaussian Noise 0.5 $\sigma\in[0.1,0.2]$
Flip Intensities 0.5 None
Sharpness 0.25 sharpness=5
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.25 contrast$\in[0.5,1.5]$
Table 9: High augmentations used to validate Tyche-IS. We focus on intensity transforms, to avoid inverting the prediction. For each image, an augmentation is sampled with probability $p$.
D.7 Few Shot Regime
We train PhiSeg on a subset of LIDC-IDRI, $\tilde{\mathcal{S}}$, and examine how this network generalizes compared to Tyche, where the context set is $\tilde{\mathcal{S}}$. We investigate four few-shot settings: 3, 5, 8 and 16. For each, we train 30 PhiSegs: 10 seeds to account for variability in our samples and three data augmentation regimes (none, light, normal). For each seed and each few-shot setting, we select the PhiSeg that does best on the validation set. Figure 24 shows that Tyche can leverage the data available much more effectively than PhiSeg, which fails to learn with so little samples.
Figure 24: Few-Shot Regime. Comparison between Tyche-TS and PhiSeg trained on the few-shot examples. PhiSeg fails to learn from very few samples both on max. Dice score and GED, compared to Tyche.
Appendix E Further Evaluation
Figure 25: Sample Diversity on Multi-Annotator Data. Since the sample diversity of the deterministic methods is 0, we do not show them here. Tyche-IS produces the most diverse samples, while Fine-Tuned SAM has very low diversity, despite varying clicks and bounding box locations. Higher is better.
Figure 26: Hungarian Matching on Multi-Annotator Data. Both Tyche-IS and UniverSeg perform well. Tyche-TS performs best. Higher is better.
E.1 Performances on other Metrics
Evaluating the quality of different predictions can be challenging especially when different annotators are available. We proposed best candidate Dice score and Generalized Energy Distance. Some also analysed sample diversity [95], and Hungarian Matching [68, 133]. Sample diversity consists in measuring the agreement between candidate predictions $\hat{\mathcal{Y}}$, rewarding most diverse sets of candidates:
$D_{SD}(\hat{\mathcal{Y}})=\mathbb{E}\left[d(\hat{p},\hat{p}^{\prime})\right],$ (17)
where $\hat{p},\hat{p}^{\prime}\sim\hat{\mathcal{Y}}$ and $d(\cdot,\cdot)$ is Dice score. One limitation of this metric is that it blindly rewards diversity without taking into account the natural ambiguity in the target. Ideally, when there is high ambiguity in the target, the samples are very diverse, and inversely, when there is low ambiguity, the segmentation candidates are not diverse.
In the context of stochastic predictions, Hungarian Matching [70] consists in matching the set of predictions with the set of annotations, so that an overall metric is minimized. We use negative Dice score. One limitation of this method is that it has to be adapted when the number of annotators does not match the number of prediction. The most widely used fix is to artificially inflate the number of predictions and the number of annotations to reach the least common multiple [68]. We use this strategy here as well.
Figures 25 and 26 show the performances for sample diversity and Hungarian Matching respectively.
Figure 27: Best candidate Dice score for the Single-Annotator Datasets. Top to Bottom: ACDC, PanDental, SCD, SpineWeb and WBC. Tyche performs well in general except for SpineWeb.
E.2 Per Dataset Results
We show for each dataset and each method the best candidate Dice score. We also show for Tyche and the different benchmarks Generalized Energy Distance, Hungarian Matching and Sample Diversity for the datasets with multiple annotations.
Best candidate Dice Score. Figure 27 shows the best candidate Dice score for the single-annotator datasets. Tyche performs well across datasets. Both versions of Tyche seem to dominate both types of benchmarks and be comparable to the upper bounds, except for one dataset: SpineWeb. We hypothesize that part of the performance drop is due to the nature of the structure to segment in SpineWeb: individual vertebra. Most of our training data contains single structures to segment.
Figure 28 shows the best candidate Dice score for the multi-annotator datasets. Similar conclusions can be drawn as for the single-annotator data. Some benchmarks are particularly sensitive to the data they are evaluated on, for instance SegGPT. This methods performs really well on the Prostate data but quite poorly on the STARE and the Hippocampus data. We assume that because SegGPT was designed for images of 448x448, our images of dimension 128x128 might affect performances.
Figure 28: Best candidate Dice Score for Multi-Annotator Datasets Top to bottom: Hippocampus, LIDC-IDRI, Prostate Task 1, Prostate Task 2 and STARE. Tyche performs well across datasets. (Higher is better.)
Generalized Energy Distance. Figure 29 shows Generalized Energy distance for the multi-rater datasets.
Figure 29: Generalized Energy Distance for Multi-Annotator Datasets. Top to bottom: Hippocampus, LIDC-IDRI, Prostate Task 1, Prostate Task 2 and STARE. Tyche performs well across datasets. (Lower is better.)
Hungarian Matching. Figure 30 shows the Hungarian Matching metric for the multi-rater datasets. We find that UniverSeg performs particularly well. We suspect that because we have to artificially duplicate our samples to compute this metric, the resulting scenario favors methods that are closer to the mean.
Figure 30: Hungarian Matching Dice for Multi-Annotator Datasets. Top to bottom: Hippocampus, LIDC-IDRI, Prostate Task 1, Prostate Task 2 and STARE. Tyche performs well across datasets. (Higher is better.)
Sample Diversity. Figure 31 shows sample diversity for the multi-annotator datasets. We only show the sample diversity for the methods outputing more than one segmentation candidate. For UniverSeg, SegGPT and SENet, the sample diversity is trivially 0.
Figure 31: Sample Diversity for Multi-Annotator Datasets. Top to bottom: Hippocampus, LIDC-IDRI, Prostate Task 1, Prostate Task 2 and STARE. Tyche performs well across datasets. We only show methods with diversity greater than 0. (Higher is better.)
E.3 Additional Visualizations
We show additional visualization for Tyche frameworks as well as all the benchmarks.
Single-Annotator Data
Visualizations for ACDC are shown Figure 32. We show two PanDental examples Figure 34 and Figure 35, one for each task. We show an example for SpineWeb Figure 33 and one for WBC Figure 37.
Figure 32: Example Prediction for ACDC.
Figure 33: Example Prediction for SpineWeb.
Figure 34: Example Prediction for PanDental, task 1.
Figure 35: Example Prediction for PanDental, task 2.
Figure 36: Example Prediction for SCD.
Figure 37: Example Prediction for WBC.
Multi-Annotator Data
We show an example prediction for the Hippocampus data Figure 38 and one example prediction for STARE in Figure 39. We provide visualizations for each Prostate task Figures 40 and 41 respectively. Finally, we give two example visualizations for LIDC-IDRI Figures 42 and 43.
Figure 38: Example Prediction for Hippocampus.
Figure 39: Example Prediction for STARE.
Figure 40: Example Prediction for Prostate Task 1.
Figure 41: Example Prediction for Prostate Task 2.
Figure 42: Example Prediction for LIDC-IDRI.
Figure 43: Example Prediction for LIDC-IDRI.
Dataset Name Description # of Scans Image Modalities
ACDC [14]
Left and right ventricular endocardium
99
cine-MRI
AMOS [54]
Abdominal organ segmentation
240
CT, MRI
BBBC003 [83]
Mouse embryos
15
Microscopy
BBBC038 [20]
Nuclei images
670
Microscopy
BUID [2]
Breast tumors
647
Ultrasound
BrainDev. [38, 37, 72, 114]
Adult and Neonatal Brain Atlases
53
multi-modal MRI
BRATS [7, 8, 93]
Brain tumors
6,096
multi-modal MRI
BTCV [75]
Abdominal Organs
30
CT
BUS [136]
Breast tumor
163
Ultrasound
CAMUS [77]
Four-chamber and Apical two-chamber heart
500
Ultrasound
CDemris [56]
Human Left Atrial Wall
60
CMR
CHAOS [60, 58]
Abdominal organs (liver, kidneys, spleen)
40
CT, T2-weighted MRI
CheXplanation [110]
Chest X-Ray observations
170
X-Ray
CT-ORG[106]
Abdominal organ segmentation (overlap with LiTS)
140
CT
DRIVE [122]
Blood vessels in retinal images
20
Optical camera
EOphtha [27]
Eye Microaneurysms and Diabetic Retinopathy
102
Optical camera
FeTA [102]
Fetal brain structures
80
Fetal MRI
FetoPlac [10]
Placenta vessel
6
Fetoscopic optical camera
HMC-QU [28, 65]
4-chamber (A4C) and apical 2-chamber (A2C) left wall
292
Ultrasound
HipXRay [40]
Ilium and femur
140
X-Ray
I2CVB [78]
Prostate (peripheral zone, central gland)
19
T2-weighted MRI
IDRID [103]
Diabetic Retinopathy
54
Optical camera
ISLES [43]
Ischemic stroke lesion
180
multi-modal MRI
KiTS [42]
Kidney and kidney tumor
210
CT
LGGFlair [18, 91]
TCIA lower-grade glioma brain tumor
110
MRI
LiTS [16]
Liver Tumor
131
CT
LUNA [115]
Lungs
888
CT
MCIC [36]
Multi-site Brain regions of Schizophrenic patients
390
T1-weighted MRI
MSD [118]
Collection of 10 Medical Segmentation Datasets
3,225
CT, multi-modal MRI
NCI-ISBI [17]
Prostate
30
T2-weighted MRI
OASIS [46, 88]
Brain anatomy
414
T1-weighted MRI
OCTA500 [79]
Retinal vascular
500
OCT/OCTA
PanDental [1]
Mandible and Teeth
215
X-Ray
PAXRay [112]
Thoracic organs
880
X-Ray
PROMISE12 [82]
Prostate
37
T2-weighted MRI
PPMI [89]
Brain regions of Parkinson patients
1,130
T1-weighted MRI
ROSE [86]
Retinal vessel
117
OCT/OCTA
SCD [104]
Sunnybrook Cardiac Multi-Dataset Collection
100
cine-MRI
SegTHOR [74]
Thoracic organs (heart, trachea, esophagus)
40
CT
SpineWeb [138]
Vertebrae
15
T2-weighted MRI
ToothSeg [50]
Individual teeth
598
X-Ray
WBC [139]
White blood cell and nucleus
400
Microscopy
WMH [71]
White matter hyper-intensities
60
multi-modal MRI
WORD [85]
Organ segmentation
120
CT
Table 10: Collection of datasets in MegaMedical 2.0. The entry number of scans is the number of unique (subject, modality) pairs for each dataset.
Dataset Name Description # of Scans Image Modalities
LIDC-IDRI [4]
Lung Nodules
1018
CT
QUBIQ [92]
Brain, kidney, pancreas and prostate
209
MRI T1, Multimodal MRI, CT
STARE [47]
Blood vessels in retinal images
20
Optical camera
Table 11: Multi-Annotator Data. The entry number of scans is the number of unique (subject, modality) pairs for each dataset.